OpenStudy (anonymous):

can anyone give me an example of integration by partial fraction?and by parts?

OpenStudy (anonymous):

Partial fractions:\[\int\limits_{a}^{b} x / (x + 1)^{2}\] Parts: \[\int\limits_{a}^{b} x \sin x dx\] \[\int\limits_{a}^{b} lnx dx\]